OpenStudy (faolan):

Simplify the following expression. x^5 x^-5 x^2

OpenStudy (faolan):

is it x^2?

Nnesha (nnesha):

when you multiply same bases you should add their exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (faolan):

what about the negative and positive 5? does that cancel those out?

Nnesha (nnesha):

yes right!

OpenStudy (faolan):

so its x^2

Nnesha (nnesha):

yes right

OpenStudy (faolan):

ok. thx!

Nnesha (nnesha):

5-5= 0 so x^0 anything to the zero power is jsut equal to one \[\rm (anything)^0= 1\]

Nnesha (nnesha):

my pleasure :)